OpenStudy (anonymous):

Solve by subsitution y= 2/3x-2 2x+3y=9

OpenStudy (anonymous):

well if y = 2/3 x - 2 then obviously we can replace y in the second equation, hence \[2x + 3(2/3 x - 2) = 9\]\[2x + 2x - 6 =9 \]\[4x = 15\]\[x = 15/4\] Substituting that into our equation: \[y = 2/3 \times 15/4 -2\]\[y= 30/12 -2 = 1/2\]\[y= 1/2, x=15/4\]